OpenStudy (anonymous):

Jamal and Tanya use two different containers to carry water to a pool. -Jamal makes A trips to the pool, and Tanya makes B trips to the pool. -Jamal’s container holds x gallons of water, and Tanya’s container holds y gallons. Write an expression that represents the total amount of water carried to the pool.

OpenStudy (faiqraees):

Total amount = Jamal trips*jamal's gallon+Tanya trips*Tanya's gallon Total amount = A*x+B*y Total amount = Ax+By